OpenStudy (icalibear):

Solve the system of equations using elimination method? 2y = x + 1 –2x – y = 7

OpenStudy (anonymous):

multiply equation 1 by two and then add with the equation two u will get y then substitute y in the equation one original...:) u will get x..

OpenStudy (anonymous):

solve it then teell me where u stuck..

OpenStudy (anonymous):

First, bring the x and y-terms to the same side and the constant to the other side. 2y = x + 1 -x + 2y = 1 Now set up a system of linear equations. -x + 2y = 1 -2x - y = 7 -x + 2y = 1 2(-2x - y = 7) -x + 2y = 1 -4x - 2y = 14 The y-terms cancel. -5x = 15 x = -3 Now, plug -3 back in for x and solve for y to get your complete solution set.

OpenStudy (icalibear):

(-3,-2)???

OpenStudy (anonymous):

yup..